kevin's sister mary is 5 years younger than he is. the total of their ages is 33 years. how old is kevin?
Answer by mananth(16946)

kevin x years
Mary x-5 years
...
x+x-5 =33
2x-5 =33
add 5 to both sides
2x=38
/2
x=19 Kevin's age
...
Mary's age = 19-5=14 years